Champion Contractors completed the following transactions involving equipment. Year 1 Jan. 1 Paid $302,000 cash plus $12,080Journal entry worksheet 1 2 3 4 5 6 Paid $302,000 cash plus $12,080 in sales tax and $1,900 in transportation (FOB shipping pPaid $4,000 to install air conditioning in the loader to enable operations under harsher conditions. This increased the estimJournal entry worksheet < 1 2 3 4 5 6 Recorded annual straight-line depreciation on the loader. Note: Enter debits before creJournal entry worksheet < 1 2 3 5 6 Paid $4,900 to overhaul the loaders engine, which increased the loaders estimated usefuJournal entry worksheet < 1 2 3 4 Paid $1,225 for minor repairs to the loader after the operator backed it into a tree. Note:Journal entry worksheet < 1 2 3 4 5 6 Recorded annual straight-line depreciation on the loader. Note: Enter debits before cre
Answer #1
Date General Journal Debit Credit
Jan 1,Year 1 Equipment 315,980 =302000+12080+1900
Cash 315,980
Jan 3,Year 1 Equipment 4,000
Cash 4,000
Dec 31,Year 1 Depreciation expense—Equipment 72,145 =(315980+4000-30200-1200)/4
Accumulated depreciation—Equipment 72,145
Jan 1,Year 2 Equipment 4,900
Cash 4,900
Feb 17,Year 2 Repairs expense—Equipment 1225
Cash 1,225
Dec 31,Year 2 Depreciation expense—Equipment 44,267 =(315980+4000+4900-30200-1200-72145)/5
Accumulated depreciation—Equipment 44,267